OpenStudy (anonymous):

When the integer N is divided by 7, the quotient is Q and the remainder is 5. When N + 24 is divided by 7, the remainder is _?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

What is the remainder when 27 is divided by 7?

OpenStudy (anonymous):

6, correct?

OpenStudy (anonymous):

No. Remainder will be 1.

OpenStudy (anonymous):

How come, Easy? Explain, please.

Directrix (directrix):

Whoops, I wrote the wrong number. 24 divided by 7 - the remainder is what?

OpenStudy (anonymous):

3 would be the remainder.

Directrix (directrix):

Here's what I am thinking: (n + 24) / 7 = n/7 + 24/7 --> Do you agree with that?

OpenStudy (anonymous):

Yep.

Directrix (directrix):

(n + 24) / 7 = n/7 + 24/7 n/7 has a remainder of 5 (Given) 24/7 has a remainder of 3 What is the sum of the two remainders? @SHSAT-Help

OpenStudy (anonymous):

5 + 3 = 8.

OpenStudy (anonymous):

Though, that's not one of the answer choices. The answer choices are: 4 3 2 1 0

Directrix (directrix):

Now, divide 8 by 7 to get the remainder when (n + 24) / 7 is divided by 7.

Directrix (directrix):

8/ 7 has what remainder?

OpenStudy (anonymous):

One.

Directrix (directrix):

Yes.

OpenStudy (anonymous):

That's the answer? o-o

Directrix (directrix):

Suppose n were 12. Then, n/7 = 12/7 which has a remainder of 5. So, what is the remainder when (12 + 24) is divided by 7? 36/7 has a remainder of what?

OpenStudy (anonymous):

One.

OpenStudy (anonymous):

So to answer these type of question, you can guess and check?

Directrix (directrix):

That is not a good technique. I gave that as an illustration of the "concept work" from earlier posts in the thread.

OpenStudy (anonymous):

Oh yeah, that reminds me. @Easyaspi314 don't give out answer without explaining. It's against the CoC, and people like me want to learn!

OpenStudy (anonymous):

Oh. Let me review your steps for a few minute.

OpenStudy (anonymous):

"Now, divide 8 by 7 to get the remainder when (n + 24) / 7 is divided by 7." I don't get that part.

OpenStudy (anonymous):

@Directrix